OpenStudy (anonymous):

Find the constant of variation for the quadratic variation. 6y = 4x2 (Points : 1) 4/9 2/3 -2 3/2

Nnesha (nnesha):

y = kx equation for direct variation where k is constant of variation so you have to solve equation for y

OpenStudy (here_to_help15):

y = (4/6)x^2 y = (2/3)x^2